4/9 of a number is 12. What is the number?
Answer by fractalier(6550)

"Of" normally means multiply...it does here...so 4/9 of a number (call it x) is 12, or
(4/9)x = 12
Now multiply both sides by 9/4 and get
x = (9/4)(12) = 108/4 = 27